OpenStudy (anonymous):

Find angle between given vectors to the nearest tenth of a degree u = <15,-9>, v = <13,-22>

OpenStudy (dumbcow):

\[\cos \theta = \frac{u*v}{|u| |v|}\] where \[u*v = (15*13)+ (-9*-22)\] \[|u| = \sqrt{15^{2} +(-9)^{2}}\] |v| in same way that should help

OpenStudy (anonymous):

thank you so much your a life saver

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

so then for v I just plug in the numbers right?